• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 18:21:08 on July 25, 2006

    Entry number 175208

    keyword=Qasy feedback turned on

    The Hall A IA slope was measured to be +1630 ppm. The Qasy feedback is enabled now.
    Some data starting at about 5:30 pm had large Qasy, but I guess this doesn't matter (e.g.
    Hall C is offline ?). The Qasy matters for G0 experiment in Hall C because the beams couple.

    I'll measure again later more accurately when beam is more "on" all the time. The slope
    will be checked once a week at least.

    See also